### v3

# LSHBloom: Internet-Scale Text Deduplication

Abstract.
[…]
PVLDB Reference Format:
PVLDB, 19(1): XXX-XXX, 2025.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
1. Introduction
Modern large language model (LLM) training pipelines require the ingestion of many millions of documents to enable internet-scale language pre-training. These documents are obtained through web scraping, deals with academic institutions, and relationships with publishers. The sheer scale of this data ingestion process presents a substantial data management and processing burden when assembling large-scale text databases for continuous LLM pre-training. LLM performance has also been shown to be sensitive to data quality (alignment of data with downstream tasks) (tan202415pintstechnicalreportpretraining) as well as data redundancy–excessive redundancy can cause performance scaling below the predicted rate of established scaling laws (chen-etal-2025-revisiting). Moreover, per-epoch training costs scale in the number of training data points. As such, it is highly desirable to remove outright uninformative data points from internet-scale text datasets–to alleviate the data management burden, to reduce the costs of pre-training, and to enhance LLM performance.
In this work, we focus on reducing data redundancy in text. Duplicate text data–that is, instances of the same content or content that only subtly differ from one another due to differences in publication format, copyediting, and versioning–are ubiquitous in internet-scale text datasets and are well known to lead to a host of undesirable outcomes, from excessive training times to the introduction of predictive biases and training data memorization (lee2021deduplicating), in trained models. Text duplication between training and evaluation sets also hinders thorough, unbiased assessments of model performance, with much evidence suggesting that standard open-source benchmarks have leaked into LLM training sets (Chandran2024PrivateBT), illegitimately inflating perceived model performance by 8–11% (Dekoninck2024EvadingDC).
These considerations motivate us to address the accurate and efficient deduplication of large numbers (100s of millions) of documents when preprocessing internet-scale text datasets.
We demonstrate via careful evaluation that existing state-of-the-art approaches to text deduplication suffer from poor deduplication performance (i.e., low precision and/or recall) and/or excessive computational costs. Due to the frequent loss of metadata and subtle variation in documents, deduplication of technical documents requires approximate textual comparison methods such as Jaccard similarity, which may be accelerated via the approximating MinHashing method (broder1998min). The infeasibility of pairwise comparisons in large datasets motivated the development of MinHashLSH (leskovec2014finding), which applies a locality-sensitive hashing (LSH) scheme to MinHash signatures to avoid comparing dissimilar documents. MinHashLSH is by far the most popular method for text deduplication in LLM data pipelines; however, the associated MinHashLSH index can grow extremely large. For example, applying MinHashLSH to the peS2o dataset (peS2o) of just 39M documents takes more than 35 hours on a single 32-core node of the Polaris supercomputer and requires more than 200 GB of disk space (see subsubsection 5.4.1).
[…]
MinHashLSH is the most popular and performant state-of-the-art text deduplication technique available today; LSHBloom can be employed as a drop-in replacement for MinHashLSH–preserving this superior performance while enabling significant gains in scaling capacity as we increase our training dataset sizes. We know from historical data that LLM training dataset sizes will almost quadruple each year (epoch2024datasetsizetrend). LSHBloom presents an opportunity to create robust data processing infrastructure for LLM pre-training that is effective, efficient, and scalable well into the future. In summary, the contributions of our work are the following:
  1. (1)
    We present the first side-by-side benchmark comparison of state-of-the-art text deduplication algorithms utilized LLM data processing pipelines, comparing both their deduplication performance (in terms of precision, recall, and F1) as well as their resource requirements as we increase the size of the underlying text data
  2. (2)
    We present the first study on the effects of tuning the hyperparameters for each of these algorithms, and present the best parameter settings for each
  3. (3)
    We develop LSHBloom, a novel text deduplication algorithm that outperforms or matches the existing state-of-the-art in deduplication performance while requiring significantly fewer resources and therefore enhancing scalability for the internet-scale data processing regime
2. Background
We first provide background information on the deduplication problem and describe the current state-of-the-art for large-scale text deduplication pipelines.
2.1. Defining the Deduplication Problem
Text deduplication algorithms seek to identify and remove duplicate instances of the same or similar content from a document, dataset, or text corpus. These methods can be applied at different resolutions (e.g., to documents, paragraphs, sentences, or even tokens) and scopes (e.g., across a corpus or within a single document). Deduplication methods may also differ in how they define duplication (e.g., exact matches, n-gram overlap, “fuzzy”). In evaluating a deduplication method for a particular application, we are concerned with its precision, recall, and computational performance. Low precision, i.e., excessive false positives, is harmful as it unnecessarily reduces the size of the training dataset and limits an LLM’s ability to learn the underlying distribution of the text data. Low recall, i.e., excessive false negatives, admits a great degree of duplicate content in our dataset, with negative effects on learning as well as potentially inducing degenerate behavior such as memorization. High computational performance is critical due to the extreme size of modern LLM datasets. For example, peS2o (peS2o) contains roughly 39 million documents, Google Scholar is estimated to index some 400 million academic papers and other scholarly literature (gusenbauer2019google), DOLMA (soldaini_dolma_2024) contains roughly 5B documents, and RedPajama (together2023redpajama) has just over 100B documents—scales at which naive pairwise comparisons become infeasible. Text deduplication algorithms are needed that can efficiently process billions of documents while maintaining high precision and recall.
Here, we focus on methods for document-level deduplication. Suppose an operator $S(d_{i},d_{j}):D\times D\rightarrow[0,1]$ that takes as input two documents and outputs a similarity score between 0 and 1. We define our task as follows: Given a corpus of $N$ documents, $D_{N}=$ $d_{1}$, …, $d_{N}$}, and a similarity threshold, $T\in[0,1]$, identify every document $d$ such that $\exists d^{\prime}\in D_{N}$ such that $d^{\prime}\neq d$ and $S(d,d^{\prime})\geq T$.
Just how to define duplication is a subtle question. For example, one could stipulate simply that two documents must be byte-for-byte identical: equivalent to setting $S$ to measure byte overlap and $T$ = 1.0 above. However, this definition may exclude many document pairs that we would like to consider duplicates due to great similarity in their token distributions or semantic content, as are encountered frequently in practical text ingestion pipelines. For example, the same document may be available from two separate venues (say, a pre-print on Arxiv and a camera-ready conference publication) that differ only slightly due to editorial decisions. When web-scraping text or obtaining articles from publishers, text data may be obtained in varied formats (e.g., HTML, PDF, XML), and the appropriate format-specific text parsing pipelines will induce small differences even when applied to the same underlying document. Even simply parsing two PDFs of the same document introduces complications—for instance, the choice of PDF parser may be influenced by the document’s origin, publisher, or language. PDF parsers that rely on optical character recognition (OCR) can also introduce artifacts that render exact matching woefully ineffective.
For these reasons, many contemporary deduplication techniques take a ”fuzzy” perspective on duplication, allowing for minor differences between items in duplicate pairs. One way to do this is to alter one’s choice of similarity metric to encapsulate a broader notion of similarity. This is common in so-called semantic deduplication techniques such as SemDeDup (abbas2023semdedup), which employ embeddings from pre-trained language models to identify semantically similar text data. Unfortunately, these techniques require training language models and performing inference on each data point before clustering similar items to identify duplicates, which is prohibitively expensive in the extreme-scale regime with billions of documents. It is more efficient to instead establish a similarity metric based on the raw text content of each document.
2.2. Jaccard Similarity and MinHash Approximation
Popular techniques such as those used to build the DOLMA (soldaini_dolma_2024) and DataComp-LM (li2024datacomp) datasets measure n-gram overlap between documents to identify duplicates. Alternatively, some techniques—such as MinHashLSH—choose a set-theoretic notion of document similarity.
In this paradigm, a text document is viewed as a bag of n-grams, and the similarity between two such sets, $A$ and $B$, defined in terms of the Jaccard similarity: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$.
However, computing Jaccard similarity requires evaluating set membership for each item present in the union of both sets. For documents, this set size can be large, and computing the Jaccard similarity explicitly can be time-consuming. The MinHashing procedure (broder1998min) provides a means of efficiently approximating the Jaccard similarity of two documents. Rather than compare the elements of each set directly, this procedure randomly permutes the order of set elements to produce so-called MinHash signatures. The probability that the MinHash signatures for two sets are equal is exactly equal to the Jaccard similarity of the two sets (broder1998min; leskovec2014finding). Therefore, we can repeatedly apply the MinHashing procedure to any two documents and compare the resulting hashes for equality to obtain an estimate of their Jaccard similarity. This estimate grows more accurate as we increase our sample size.
2.3. MinHashLSH for Scalable MinHash
The use of MinHash-estimated Jaccard similarity, $J_{H}$, as an approximation for Jaccard similarity, $J$, provides for more efficient similarity computations because the procedure employed to estimate the similarity of two documents has linear complexity in the number of random permutations applied, rather than quadratic complexity in the number of n-grams present in each document. However, the need to compare every pair of documents for similarity during deduplication still introduces quadratic cost-scaling with respect to the number of documents in our dataset.
Fortunately, there exists a technique that allows us to compare only those pairs with a high likelihood of similarity, rather than all pairs. MinHashLSH (leskovec2014finding) applies a locality-sensitive hashing scheme for MinHash signatures to avoid comparing dissimilar documents. Given a set of MinHash signatures, one can construct a MinHash signature matrix in which each column contains all MinHash values for a single document. Then, given a desired Jaccard similarity threshold, $T$, one can determine an optimal band size, $b$, with which to group the rows of the MinHash signature matrix. This results in $b$ groupings of rows (bands) of size $r$. With high probability, if two bands of the MinHash signature matrix are exactly identical for two documents (columns), then those two documents have Jaccard similarity greater than or equal to $T$. This procedure enables us to produce a series of candidate pairs of documents that we can take as duplicate pairs or further evaluate for similarity. This method avoids pairwise comparison between every pair of documents in the corpus but introduces associated false positive and false negative probabilities due to the use of locality-sensitive hashing.
One can implement a prefix-tree or suffix array to search efficiently for candidate duplicate documents when evaluating a new document for duplication. However, this requires storing a representation of each document, and while MinHash representations are small compared to raw documents, the MinHashLSH index becomes large for large datasets: e.g., 277 TB for 5B documents. Additionally, inserting and querying these tree-structured indices constitutes a significant portion (¿85%) of algorithm runtime, which can become prohibitive at extreme scales (see Figure 1). Here we remedy this space and runtime issue by employing a novel combination of MinHashLSH and Bloom filters.
Figure 1. Breakdown of wall clock time on 10% of peS2o for the conventional MinHashLSH algorithm and our LSHBloom method
2.4. Bloom Filter
The Bloom filter (bloom1970space) is a space-efficient probabilistic data structure used to test whether an element is a member of a set. It can determine with certainty if an element is not in the set, but if it indicates the element is in the set, there is a small probability of error. That is, the Bloom filter may admit false positives but not false negatives.
A Bloom filter consists of a bit array of size $m$, with all bits initially set to 0. $k$ independent hash functions are defined, each of which hashes an element to an integer in the range 1 to $m$. To insert an element into the Bloom filter, it is passed through the $k$ hash functions, and the resulting $k$ positions in the bit array are set to 1. We can then check if an element is in the set by hashing using the same $k$ hash functions. If all $k$ of the corresponding positions in the bit array are 1, the filter returns a positive. If any of those $k$ bits is 0, the element is not contained in the set.
Bloom filters have been used extensively in distributed computing for a variety of purposes (broder2004network; chervenak2008globus; tarkoma2011theory).
3. Related Work
Having provided background on efficient document similarity measures, we briefly review related work on deduplication methods employed by groups developing LLMs. LLM developers use various methods, many ad hoc, to deduplicate the large quantities of text content used for LLM training regardless of whether the text comes from a traditional source or synthetic data. Many leading LLMs do not discuss the deduplication and curation methods they use, such as Llama-2 (touvron_llama_2023-1), Claude, Gemini (gemini_team_gemini_2024; gemini_team_gemini_2024-1), and GPT versions 3 and later (openai_gpt-4_2024). Others, such as Palm2 (anil2023palm) and Gemma (gemma_team_gemma_2024; gemini_team_gemini_2024-1), provide only limited details. Others, including several leading models, provide extensive details about how data for their model was deduplicated including Olmo’s Dolma dataset (soldaini_dolma_2024) as well as its successor DataComp-LM (li2024datacomp), Llama 1 (touvron_llama_2023), Llama 3 (dubey_llama_2024), and OPT (zhang_opt_2022). For example, Llama 3 (dubey_llama_2024) was trained on a dataset that was deduplicated by using MinHash over the entire dataset, line-level deduplication using CCNet for lines replicated over six times in each bucket of text based on coarse categorization (e.g., math reasoning, geometry/trigonometry), and URL deduplication for web data. Phi uses an approach described by gunasekar_textbooks_2023 and elaborated on in their Phi 2.0 report (javaheripi_surprising_2023) that starts from deduplicated code and text, but they do not discuss deduplicating after generation. C4 (t5_c4_raffel2020) applies only heuristic filtering for data quality (e.g., removing lines without punctuation or fewer than three words). RedPajama (together2023redpajama) uses both exact matching via Bloom filters and fuzzy deduplication via MinHashLSH. Datasets such as the Pile (gao2020pile800gbdatasetdiverse) and more recent releases such as FineWeb2 (penedo2025fineweb2pipelinescale) apply MinHashLSH for fuzzy deduplication.
The most common methods used across these model sources include metadata-based approaches, MinHashLSH, CCNet, and n-gram-based methods. Metadata-based approaches (e.g., exact URL matching, DOIs) are by far the simplest approach; however, they are applicable only in cases in which a reliable source of metadata exists, such as those provided by a single publisher. Unfortunately, for our target datasets this metadata is often unavailable or is otherwise untrustworthy. This can occur, for example, when: 1) a document appears in two or more collections with different metadata schemes (e.g., ArXiV vs. ACM, different GitHub repositories); 2) sections/paragraphs of text are copied, as in extended journal versions of computer science papers; or 3) metadata is not curated (e.g., an ArXiV entry may not be updated to reflect a subsequent publication).
CCNet (wenzek2020ccnet) (ccnet) is an automated preprocessing pipeline that aims to extract high-quality monolingual language datasets from the Common Crawl corpus of Web content. CCNet performs dataset sharding, deduplication, language identification, and quality filtering. CCNet’s deduplication technique begins by preprocessing text so that all characters are lowercase and eliminates special Unicode characters. CCNet then splits documents by newline characters, computes SHA1 hashes for each resulting unit of text, and compares SHA1 hashes between units of text to identify duplicates. CCNet can optionally compare to a single document, a subset of documents, or all documents.
Dolma (soldaini_dolma_2024) deduplicates paragraphs by exact matching using a Bloom filter (dolma). An alternative procedure (dolma-ngram) splits paragraphs into n-grams and checks whether n-grams are duplicated by querying against a Bloom filter. If the proportion of duplicated n-grams in a paragraph exceeds a threshold, $T\in$ [0, 1], then that paragraph is marked as a duplicate. By default, $T$ = 1.0 in their configuration.
DataComp-LM (li2024datacomp) (dclm) deduplicates items at the document and paragraph levels simultaneously by using a Big Friendly Filter, an extension of the Bloom filter. They split each document into paragraphs on newline characters and tokenize paragraphs with the UniSeg tokenizer (ye2023uniseg). For each paragraph, they query each n-gram against their Bloom filter and remove duplicated n-grams. If the proportion of duplicated n-grams in a paragraph exceeds a user-defined threshold, $T\in$ [0, 1], the entire paragraph is removed from the document. If the proportion of duplicated n-grams in the entire document exceeds the user-defined threshold, then the document is removed from the dataset. They employ the same threshold parameters for paragraph-level and document-level deduplication.
A table comparing various methods for data deduplication in contemporary LLM pre-training pipelines can be found in Appendix A.
4. The LSHBloom Method
4.1. Overview
We propose a memory-efficient alternative to MinHashLSH for document deduplication. In particular, rather than employ the traditional prefix-tree index used to find exact matches between the MinHash signatures in each band, we develop a novel Bloom filter index structure for MinHashLSH. We instantiate a series of Bloom filters, one corresponding to each band of the MinHash signature matrix. By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for duplicates. The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of bender2022optimal.
4.2. Insertion
Given a Jaccard similarity threshold $T\in$ [0, 1], MinHashLSH traditionally groups signatures into a series of $b$ bands of size $r$ such that an exact match in one band for two columns of the MinHash signature matrix denotes a duplicate document. That is, with high probability, the two documents with a matching band have Jaccard similarity greater than or equal to the threshold. To avoid storing the signatures in full, as is required by MinHashLSH, we assign each band of the minhash signature matrix its own Bloom filter. Thus we have in total $b$ Bloom filters, and a collision in any one indicates a duplicate document.
When we wish to insert a document into the index, we first compute the MinHash signature for a document, which is a vector of integers (each produced by a universal hash function). Then, we group the rows into $b$ bands of size $r$, as dictated by the MinHashLSH algorithm. It is desirable to hash this list of integers into a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions. To do so, we take advantage of the fact that each MinHash signature value ($h_{i}(x_{i})$) is computed using a universal hash function (in our case, SHA1 (sha1)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values. Then, we can hash a vector $\bar{x}$ of $r$ integers as follows (carter1977universal):
$h(\bar{x})=\left(\sum_{i=1}^{r}h_{i}(x_{i})\right)\mathrm{mod}\ {N}$
This approach enables us to map each band of MinHash signatures to a single integer value with known collision probability ($1/N$). To insert a document into the index, we simply insert each of the $b$ final integer values into each of the $b$ Bloom filters corresponding to the $b$ bands in the signature matrix.
4.3. Querying
Querying a document against the index is a similar process to insertion. We compute the document’s MinHash signature, group the rows into bands, and then hash each band as described in subsection 4.2. If there is a collision for any band of signatures, we mark that document as a duplicate. This approach mimics the original MinHashLSH procedure and gives the same probabilistic guarantees with respect to the Jaccard similarity of candidate pairs, with a slight increase in false positive probability due to our use of Bloom filters. However, the false positive probability is configurable and can be set to be arbitrarily low.
4.4. Error Rate
In traditional MinHashLSH, a collision in any band causes a document to be marked as a candidate duplicate. This is also the case in LSHBloom. However, we make use of Bloom filters, which incur additional false positive overhead. Here, we analyze the error rate of LSHBloom. A false positive is a non-duplicate document that is incorrectly identified as a duplicate by LSHBloom, and a false negative is a duplicate document that is not marked as a duplicate by LSHBloom.
Given a Minhash signature matrix with $b$ bands with $r$ rows each corresponding to $b$ Bloom filters with their individual false positive rates set to $p$, the net false positive overhead of using Bloom filters will be $p\textsubscript{effective}=1-(1-p)^{b}$. This is because a single collision in any band constitutes a duplicate, and false positive rates are independent for each Bloom filter. Therefore, the probability that no Bloom filter reports a false positive is $(1-p)^{b}$, and the overall false positive rate overhead introduced by using Bloom filters is the complement of this value.
The relationship between the false positive rate of the Bloom filter and the overall error rate of LSHBloom can be understood as follows. leskovec2014finding determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
(1) $FP_{lsh}=\int_{0}^{T}1-(1-t^{r})^{b}dt$
(2) $FN_{lsh}=\int_{T}^{1}1-(1-(1-t^{r})^{b})dt$
zhu_lshensemble demonstrate how to tune the values of $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
A Bloom filter will not yield false negatives, but may yield false positives. If MinHashLSH yields a false positive (by “mistakenly” yielding two identical bands), we get a false positive result. If MinHashLSH does not yield a false positive, the bloom filter index may still yield a false positive result with probability $p_{effective}$. Additionally, there is a small probability of hash collision ($1/N$) when reducing each band to a single integer (as described in 4.2) for insertion into the Bloom filter. Therefore, the overall false positive rate for LSHBloom is given by:
(3) $FP_{bloom}=FP_{lsh}+(1-FP_{lsh})*(p_{effective}+b/N)$
Note that $p_{effective}$ can be set arbitrarily small in practice so that this overhead is negligible. For 32-bit hashvalues (as is the default in packages such as datasketch) $N=2^{32}$, so the probability of collision for bands prior to insertion into the Bloom filter is also quite small. Additionally, because the size of the Bloom filter index scales with the number of items inserted and is not influenced by the size of the integers inserted into it, one can increase the size of the hashvalues (to say, a 64 or 128-bit integer) without increasing the size of the Bloom filter index. This is not true of the traditional MinHashLSH index which must store these hashvalues for comparison.
Meanwhile, for each false negative that MinHashLSH produces, $p_{effective}$ percent of those false negatives will be marked as positives due to false positive hits in the bloom filter index and $b/N$ percent of them will be marked positive due to collisions when reducing the bands of the MinHash signature matrix for insertion. That is,
(4) $FN_{bloom}=(1-(p_{effective}+b/N))*FN_{lsh}$
4.5. Advantages
We seek to measure duplication based on the Jaccard similarity between set-theoretic representations of documents as bags of n-grams. Since many practical scenarios may admit an extremely large universe of such n-grams (e.g., if we allow our set of documents to include a variety of encodings, languages, latex, tablature, figures, etc.), practitioners may want to reflect this additional complexity by increasing the size of MinHash hashvalues (to 64 or 128-bits or more) or increasing the accuracy of MinHash estimates of the Jaccard similarity by increasing the number of MinHash permutations. A scalable deduplication algorithm should be amenable to such considerations.
LSHBloom allows for the space-efficient representation of extremely large deduplication indices. This is due to the Bloom filter’s ability to determine the set membership properties of objects of arbitrary size with a fixed number of bits. This means that increasing the size of hashvalues does not change the index size. Furthermore, using more MinHash permutations only increases the index size linearly with the number of bands that are required for the MinHash signature matrix–which is far fewer than the number of permutations. If we expect $n$ documents and desire a false positive probability of $p$, then each Bloom filter requires exactly $m=-n*log(p)/(log(2))^{2}$ bits (bender2022optimal).
As an example, say we use a Jaccard similarity threshold $T$ of 0.8, and 128 random permutations for MinHashing. For these parameter settings, MinHashLSH creates nine bands (using the method described in (zhu_lshensemble) to tune $b$ and $r$) for the MinHash signature matrix, meaning that we require nine Bloom filters for our index. For these settings, setting a conservative ppeffective = 1e-10 for a dataset of 10B documents would only require 590 GB of space. Meanwhile, traditional MinHashLSH would require 46 TB (nearly $80\times$ more space) to store the equivalent LSHIndex—exemplifying the scalability of our technique.
While the use of Bloom filters to satisfy our memory and runtime constraints introduces additional false positive errors, it also provides another level of configuration. This is another advantage of our technique: we can observe massive gains in space while only introducing a single additional hyperparameter, meaning that our hyperparameter tuning burden is comparable to that of MinHashLSH.
The additional false positive error overhead scales with the number of documents we intend to insert and scales inversely with the size of each filter. Therefore, there is a trade-off between memory usage and correctness that is mediated by the size of each filter. However, even for very large datasets of, say, 100M documents, this does not prove to be an issue as we can set an effective false positive probability overhead as low as 1e-15 and only incur a space cost of 8.5 GB for our index compared to 460 GB for the equivalent LSHIndex.
5. Evaluation
Broadly our evaluation considers two concerns important for large-scale deduplication: correctness and performance. As expressed above, both the correctness and performance of these methods are critical for their application to extremely large datasets. To assess correctness, we tune LSHBloom and other state-of-the-art deduplication methods on a balanced dataset of 25K PDFs containing labeled duplicates. We then use the best parameter settings for each algorithm and compare them on a larger labelled dataset of 50K PDFs. Without sufficient correctness, the system risks either under-identifying duplicates, leading to decreased downstream model quality and unnecessary training overhead; or over-identifying duplicates, which unnecessarily reduces the training dataset volume. For performance, we consider the storage utilization and runtime of each method on a larger collection of 39M PDFs, peS2o (peS2o). Without sufficient performance, it may not be possible to process large datasets on available hardware resources.
In the following subsections, we describe the details of our evaluation beginning with the evaluation methodology, then proceeding to correctness, performance, and finally the tradeoffs between performance and correctness.
5.1. Experimental Methodology
In this subsection, we describe our experimental methodology including testbed hardware, state-of-the-art methods for comparison, our evaluation metrics, and datasets.
5.1.1. Testbed Hardware
We ran our experiments on the Polaris supercomputer, a 560-node HPE Apollo 6500 Gen 10+-based system. Each node is equipped with a 32-core AMD Zen 3 2.8 GHz processor, with 512 GB of DDR4 RAM, and 3.2 TB of local SSD: representative of nodes found in large-scale high-performance computing facilities used for deduplication tasks.
5.1.2. Baselines
We compare LSHBloom with state-of-the-art deduplication methods widely used in large-scale LLM training: MinHashLSH, Dolma, Dolma-ngram, CCNet, and DataComp-LM. We normalized all implementations to Python-only implementations to avoid performance differences due to implementation language. This also enabled us to normalize the Bloom filter implementation for those techniques which use them. Below, we describe any changes or configurations to the methods.
MinhashLSH (broder1997resemblance) (lsh): We configure the traditional MinhashLSH algorithm as described in subsection 2.3. DOLMA: We apply both standard dolma and dolma-ngram methods to document-level deduplication. Specifically, for dolma we determine whether a document is duplicated by identifying duplicated paragraphs and comparing the percentage of document text duplicated to the overlap threshold, $T$. For dolma-ngram, we split each document into a set of n-grams with a given size, determine how many n-grams are duplicated according to the Bloom filter, and compare the percentage of n-grams that are duplicated within that document to the overlap threshold, $T$. Our implementation aims to faithfully extend Dolma’s two primary deduplication methods to document-level deduplication. CCNet originally performs deduplication at the paragraph level; however, here, we apply it at the document level by measuring the proportion of duplicated paragraphs in any given document against some tolerance threshold, $T\in$ [0, 1]. For a fair comparison, we compare documents to all other documents—just as MinhashLSH does. When evaluating memory usage, we apply all CCNet optimizations, such as the use of memory-efficient hash sets. Additionally, we augment CCNet with multiprocessing where appropriate (e.g., when computing hashes over each document) for a more appropriate comparison during our scaling tests. DataComp-LM: For our evaluation, we compare only with their document-level deduplication procedure.
When considering the correctness of techniques that operate on n-grams and use Bloom filters, we must give an expected n-gram count in the corpus; choosing this count accurately is crucial to make a fair comparison. Because counting the number of n-grams of a given size in a corpus is far more computationally expensive than counting the number of documents (as is required for our technique), we compute an estimate of the n-gram count in a corpus. To do so we randomly sample $N$ = 1000 documents from the corpus, compute the mean n-gram count from our sample, and multiply the mean n-gram count by the number of total documents in that corpus to obtain an estimate of the number of n-grams in the entire corpus. We parameterize Bloom filters for n-gram techniques using this estimate. We apply an analogous procedure to obtain estimated paragraph counts for Dolma and CCNet.
5.1.3. Metrics
Correctness: We require methods that can distinguish meaningfully between duplicate and non-duplicate documents. Here, we consider a ”positive” label to mean that a document is a duplicate of another document in the corpus. We want these methods to yield a few false positives (to retrain the maximal number of training examples), few false negatives to reduce the prevalence of duplication in our dataset, and to provide a reasonable combination of accuracy and performance. Particularly for large-scale deep learning, retaining the maximal number of training examples is highly desirable—therefore, we want few false positives. Similarly, allowing a large number of duplicates into the training corpus is unacceptable as this affects the model’s training efficiency and may increase the likelihood that our model develops degenerate behaviors such as memorization. To account for these considerations, we evaluate each deduplication technique based on its ability to mitigate both false positives and false negatives using its precision (the proportion of duplicate labels that are true duplicates), its recall (the proportion of duplicates in the dataset that we correctly identify), and its F1 score (the harmonic mean of precision and recall):
$F1=\frac{TP}{TP+\frac{1}{2}(FP+FN)}$
Performance: To assess the resource costs of each technique we measure wall clock time and disk space usage. We measure wall clock time using /user/bin/time and assess each method’s disk space requirements using Python’s standard os.path.getsize method.
5.1.4. Constructing the correctness evaluation dataset
To permit the rigorous evaluation of de-duplication methods, we created a pipeline for generating labeled datasets of duplicate documents. We acquire scientific articles (scraped from HTML and parsed from PDF) from a variety of sources, spanning a diverse set of disciplines (mathematics, physics, chemistry, biology, economics, engineering, and medicine) as identified through keywords or classification algorithms based on metadata. Text content from HTML is directly extracted. We translated each PDF into a sequence of tokens by applying a text extraction or OCR tool—one of PyMupDF (PyMuPDF), Nougat (blecher2023nougat), or Tesseract (smith2007overview), each used with roughly the same frequency. We then can construct datasets where documents are duplicated, either by including multiple versions of the same document which were parsed using different text extraction tools and/or by including randomly truncated versions of that document. This reflects common forms of duplication that occur in real-world web scraping pipelines where the same document can be ingested via different text extraction tooling either due to differences in the origin of the document instance, the fact that the same document is often found in multiple formats (HTML, XML, PDF, etc.), and due to parsing errors which can abruptly skip or truncate portions of the document’s text.
We build a balanced dataset of 24,000 documents for use in tuning hyperparameters for each deduplication algorithm. We further evaluate each technique using its best hyperparameter settings on datasets of 50,000 documents with increasing levels of duplication, from 10% to 90% duplicate instances.
5.1.5. Choice of values for Tuning Parameters
We evaluate each technique over a range of evenly spaced parameter settings using a grid search. Specifically, for any threshold parameters such as the Jaccard similarity threshold for MinHashLSH and the text overlap threshold for DataComp-LM, we evaluate values between 0.2 and 1.0 in steps of 0.2. We vary the number of permutations for MinHashLSH from 32 to 256 by powers of two. For n-gram techniques, we explore n-gram sizes 1, 2, 5, 7, 13, and 26. For MinHashLSH and LSHBloom, we additionally gathered results for $T=0.5$ and 48 random permutations as the surrounding parameter values demonstrated sharp jumps in the F1 score, and we wanted to explore this change in the score at a finer granularity. All Bloom filters (for DOLMA/DOLMA-ngram, DataComp-LM, LSHBloom) are assigned a false positive rate $p$ of 1e-5. For LSHBloom, this equates to $p\textsubscript{effective}$ between 1e-5 and 8e-4, depending on the value of $T$ and the number of MinHash permutations. The value of $p_{effective}$ at various parameter settings for this benchmark is illustrated in Figure 2.
Figure 2. Value of $p_{effective}$ (the effective false positive rate overhead) in our correctness benchmark at various parameter settings given NN=24,956 documents and pp=1e-5
It is possible that one could properly tune these deduplication methods further using more incremental changes in their parameter settings. However, the expensive runtime of n-gram techniques on large corpora coupled with the need to comb over parameter settings in small increments points to the inherent difficulty and cost associated with properly tuning these methods for maximal deduplication performance. Meanwhile, MinHashLSH-based techniques readily admit parameter tuning with relatively low runtime and memory costs. Furthermore, MinHashLSH and LSHBloom both benefit from established theoretical analysis that effectively bounds their false positive and false negative rates given the values for the Jaccard similarity threshold and number of permutations. MinHashLSH-based techniques offer the greatest opportunity to tune parameters for the best deduplication performance, particularly when faced with a constrained parameter tuning budget. This is particularly true for LSHBloom, which is significantly cheaper to run than MinHashLSH in terms of wall clock time and disk usage—particularly as we increase the size of our dataset to several million or several billion documents (see subsubsection 5.4.1).
(a) MinHashLSH
(b) LSHBloom
Figure 3. F1 score for LSH techniques as a function of the number of permutations (x-axis) and Jaccard similarity threshold (y axis).
5.2. Hyperparameter Tuning Results (24K Dataset)
5.2.1. LSH Methods
The parametrization of methods such as MinHashLSH and LSHBloom has dramatic effects for their accuracy so tuning the parameters is a critical aspect of deploying them for large-scale deduplication efforts. Figures 3(a) and 3(b) show the F1 score for MinHashLSH and LSHBloom, respectively, with different settings for the Jaccard similarity threshold, $T$, and the number of permutation functions for MinHashing on the 24K dataset. As the number of permutation functions increases, the MinHash estimate of the Jaccard similarity becomes more accurate, and performance generally improves, which is consistent with the theory of MinHashing (broder1998min). Correspondingly, we use 256 MinHash permutations in our subsequent experiments. The threshold parameter affects the F1 score the most, as it defines the sensitivity of the deduplication algorithm to overlap in text content. For our particular benchmark, a threshold of 0.5 yields the best performance; more stringent thresholds lead to a diminishing F1 score.
(a) DOLMA
(b) DataComp-LM
Figure 4. F1 Score for N-Gram techniques as a function of N-Gram size (x axis) and overlap threshold (y axis).
5.2.2. Paragraph Level and N-Gram Methods
Figure 5 shows the impact of adjusting the overlap threshold parameter for paragraph-level methods, DOLMA and CCNet, on the F1 score. Distinguishing documents at the granularity of individual paragraphs is highly error-prone and results in poor performance. The best F1 for paragraph-level methods is achieved by setting a low threshold of 0.2. N-gram methods such as DCLM do comparably better in terms of F1 score but still lag just behind the scores of the MinHashLSH-based techniques. Interestingly, DOLMA and DOLMA-ngram perform very similarly. Figures 4(a) and 4 show the results of varying both the n-gram size and overlap threshold parameters for each n-gram method. While n-gram methods enable a finer-grained comparison between individual documents and the rest of the corpus, these methods evaluate overlap by determining the portion of n-grams in a document that is exactly duplicated in the corpus. This is similar to the notion of evaluating the set overlap between documents, but can be sensitive to the relative frequency of n-grams within a document, resulting in increased error. Moreover, paragraph-level techniques demonstrate a lack of sensitivity to changes in their parameter settings in this benchmark, indicating bias in their predictions. Specifically, we find that paragraph-level methods tend to be biased toward marking documents as duplicates.
Table 1 shows the parameter settings for each technique that we use for the remainder of our experiments.
Figure 5. F1 score vs. threshold for paragraph-level techniques.
Table 1. Best Settings for Each Deduplication Technique
Technique N-gram Size Threshold
MinHashLSH 1 0.5
LSHBloom 1 0.5
Dolma-Ngram 5 0.2
DCLM 5 0.2
Dolma - 0.2
CCNet - 0.2
5.3. Deduplication Correctness (50K Dataset)
(a) Precision
(b) Recall
(c) F1
Figure 6. Correctness metrics for deduplication techniques for various levels of duplication.
5.3.1. Deduplication Quality
Here we evaluate the correctness of each technique using metrics such as their precision, recall, and F1 score. We do so by constructing datasets of 50,000 documents each with varying levels of duplication between 10% and 90%. We construct these datasets using the procedure described in subsubsection 5.1.4.
Figure 6 illustrates the precision, recall, and F1 score of each technique as we vary the proportion of duplicate documents in our dataset. Notably, in terms of F1 score (see 6(c)), both MinHashLSH and LSHBloom outperform all other methods—only being slightly outperformed by DCLM and Dolma-Ngram (whose performance tracks closely together) when the vast majority ($60\%$) of the dataset consists of duplicates, which is highly unlikely in real-world scenarios. MinHashLSH and LSHBloom also achieve nearly identical F1 scores across experimental settings, exemplifying that our technique can closely match the performance of MinHashLSH while affording considerable gains in scalability (to be shown below in subsection 5.4).
Looking more closely, 6(a) illustrates that MinHashLSH and LSHBloom tend to consistently outperform the other techniques in terms of precision, which is particularly noteworthy for those settings where duplicates are rare, such as when duplicates only comprise 10 or 20% of the dataset. The next best technique in terms of precision is Dolma’s paragraph-level deduplication technique, with its precision often tracking closely with that of the MinHash-based techniques. However, Dolma’s dismal recall scores (<0.4<0.4) suggest that this is likely the result of under-identifying duplicate instances rather than being a reflection of greater capacity for discernment. DCLM, Dolma-Ngram, and CCNet tend to underperform the other techniques in terms of precision. Note that there is a natural tendency for precision scores to increase as the proportion of duplicates in the dataset increases. This is due to a greater number of true positives being present in the dataset.
In terms of recall, 6(b) demonstrates that DCLM and Dolma-Ngram tend to perform the best—though notably their relatively stable recall scores come at the cost of decreased precision, particularly when duplicates are rare. This indicates that, while these techniques identify most duplicates, they struggle to differentiate between duplicate and non-duplicate entities. Meanwhile, MinHashLSH and LSHBloom provide virtually identical recall scores to one another between 0.7 and 0.9 while simultaneously demonstrating strong precision. Paragraph-level techniques demonstrate lower recall than all other techniques by a wide margin.
Our experimental results indicate that LSHBloom achieves deduplication performance comparable to the state-of-the-art—indeed, LSHBloom achieves essentially identical performance to MinHashLSH, one of the most popular choices for text deduplication, with only a marginal degradation in precision.
5.3.2. Performance vs. Quality Tradeoffs
Here we investigate tradeoffs between the costs of various deduplication techniques (in terms of runtime and disk usage) and their deduplication performance. We visualize data from our evaluation of the balanced dataset of 50,000 documents.
Figures 7(a) and 7(b) plot the F1 score for each deduplication technique against runtime and disk usage, respectively. MinHashLSH and LSHBloom dominate the alternative techniques, achieving the highest F1 scores while also exhibiting competitive runtime. LSHBloom runs faster than MinHashLSH and this runtime advantage becomes more exaggerated at scale (see subsection 5.4). LSHBloom also displays a comparable F1 score to traditional MinHashLSH while using just a fraction of the disk space for its index. As shown below in subsection 5.4, this is a crucial distinction, as rapidly increasing disk usage renders deduplication efforts infeasible at the scale of many millions or billions of documents.
Our results indicate that LSHBloom intelligently discerns between duplicate and non-duplicate documents (as demonstrated by its high precision scores) and can adeptly identify the majority of duplicate documents (as demonstrated by its strong recall scores). Our results also illustrate that the main difference in performance between MinHashLSH and LSHBloom comes in the form of slightly reduced precision for the latter. Intuitively, this makes sense as LSHBloom yields slightly more false positive predictions due to its use of the Bloom filter. Overall, however, both yield similar F1 scores. This, combined with LSHBloom’s greatly reduced resource usage compared to MinHashLSH, positions it as the best choice for text deduplication use-cases.
5.4. Deduplication Resource Usage at Scale
(a) Quality vs. Wall Clock Time
(b) Quality vs. Disk
Figure 7. Pareto plots for F1 score vs. resource usage.
As LLM training datasets grow ever larger, it is important that text deduplication strategies can scale effectively to larger quantities of documents. Here we work with the peS2o dataset (peS2o) of roughly 39M documents. We evaluate each deduplication method on subsets of peS2o of various sizes and compare them in terms of wall clock time and disk usage (e.g., for storing an index or Bloom filter). We do not evaluate the n-gram techniques for scaling capacity on peS2o due to their prohibitive cost. DataComp-LM, for example, requires almost three hours just to deduplicate 50,000 examples in our previous benchmark in subsection 5.3. Even processing just 10% of peS2o using this technique would require over 200 hours by extrapolation. We discuss the results for other methods below.
Figure 8. Disk size of index for different methods, as measured for peS2o subsets of increasing size.
Figure 9. Wall clock times for different methods, as measured for peS2o subsets of increasing size.
5.4.1. Scaling of Resource Usage
In this section, we study how resource usage grows with the number of documents. In these experiments, we use the same settings shown in Table 1. We set the effective false positive rate for all bloom filters to 1e-10.
[…]
5.4.2. Estimated Resource Usage at Scale
Figure 10. Predicted wall clock times vs. number of documents (in Billions) for different methods
[…]
Table 2. Predicted index sizes for 5B documents, for LSH and LSHBloom with different false positive overheads.
Technique Bloom filter false Index disk
positive overhead usage (TB)
MinHashLSH 277.68
LSHBloom 1e-5 8.33
LSHBloom 1e-8 12.11
LSHBloom 1/5B 15.50
Table 3. Predicted index sizes for 100B documents, for LSH and LSHBloom with different false positive overheads.
Technique Bloom filter false Index disk
positive overhead usage (TB)
MinHashLSH 555.35
LSHBloom 1e-5 16.66
LSHBloom 1e-8 24.21
LSHBloom 1/100B 31.76
6. Conclusions
[…]
In the future, we aim to further reduce the wall clock time of LSHBloom by employing parallelization over subsets of text datasets when inserting them into our index—an approach that we expect to enable significant acceleration. Additionally, though very useful, our synthetic benchmark has drawbacks—namely, in its biased and limited notion of what constitutes a proper duplicate. Ideally, future work will involve benchmark datasets that cover more diverse forms of duplication in proportion to their representation ”in the wild”—that is, in real-world LLM pre-training data pipelines.
Appendix A Comparison of Data Deduplication Pipelines for LLM Pre-training
Table 4. Comparative Analysis of Data Deduplication Pipelines for LLM Pre-training
Dataset / Model Primary Data Sources Key Cleaning & Filtering Steps Dedup. Level Deduplication Technique(s) Key Parameters & Implementation Notes Citations
C4 (for T5) Common Crawl (April 2019 snapshot) Heuristic filtering: remove lines without terminal punctuation, lines << 3 words, docs << 5 sentences, docs with ”Lorem ipsum”, and docs containing words from a blocklist. Language filtering for English (>>99% probability). None Heuristic Filtering Only No explicit deduplication algorithms were used. The focus was on rule-based filtering. They remove all but one instance of any three-sentence span that occurs more than once in the dataset. (t5_c4_raffel2020)
The Pile (for GPT-Neo) 22 diverse sources including Common Crawl (Pile-CC), Books3, arXiv, GitHub, OpenWebText2, and PubMed Central. Source-specific cleaning (e.g., converting TeX to Markdown for arXiv, parsing top answers for StackExchange). The entire dataset was filtered for duplicates. Document MinHashLSH The OpenWebText2 component specifically used in-memory MinHashLSH via the datasketch library for near-duplicate detection. (gao2020pile800gbdatasetdiverse)
GPT-3 A mix of Common Crawl (filtered), WebText2, Books1, Books2, and Wikipedia. Quality filtering was applied to the Common Crawl portion. Document MinHashLSH Spark’s MinHashLSH implementation with 10 hashes. (brown2020language)
RedPajama / SlimPajama (LLaMA Replication) LLaMA sources: Common Crawl, C4, GitHub, Books, arXiv, StackExchange. Multi-stage pipeline: NFC Unicode normalization, filtering documents << 200 characters. Document Exact Hashing (Bloom Filter) & Fuzzy Deduplication (MinHashLSH) RedPajama-V2 uses a Bloom filter for efficient exact deduplication in addition to MinHashLSH. SlimPajama uses a distributed MinHashLSH implementation with 13-grams. (together2023redpajama; cerebras2023slimpajama)
CCNet CommonCrawl Language identification (for splitting into monolingual datasets) using lightweight classifiers. LM-based quality filtering at the document level. Exact (SHA1 Hashing) Line-level Computes SHA1 hashes for each line of text and performs exact deduplication at the line-level using these hash values. (wenzek2020ccnet)
Llama 3 Constructed ”from a variety of data sources containing knowledge until the end of 2023”. PII and safety filtering. Heuristic filtering: removed lines with repeated content, removed lines with ”dirty” words, filtered out documents with excessive numbers of outlier tokens compared to the rest of the training corpus. Document and Line-Level URL (metadata), MinHash, CCNet The paper mentions the use of URL deduplication, MinHash for document-level deduplication, and CCNet for line-level deduplication. (dubey_llama_2024)
Dolma Collected from CommonCrawl, Github, Reddit, Wikipedia, among other sources. Remove data containing PII or toxic material. Not Specified Paragraph Deduplicates paragraphs via exact-matching (Bloom filter). Alternatively, deduplicates paragraphs by measuring their n-gram overlap with the rest of the corpus, also using a Bloom filter. Preprocess some web data using CCNet and metadata-based deduplication. (soldaini_dolma_2024)
DataComp-LM (DCLM) Common Crawl (provides a candidate pool of 240T+ tokens). A benchmark for data curation strategies. The DCLM-Baseline dataset was created using model-based filtering (data quality classifiers) to select high-quality data. Paragraph and Document-level Baseline dataset is deduplicated using Bloom filters to measure n-gram overlap of a paragraph/document with the rest of the dataset. The DCLM-Baseline dataset of ~4T tokens is combined with StarCoder and ProofPile2. (li2024datacomp)
FineWeb / FineWeb-Edu / FineWeb2 Common Crawl (96 snapshots for FineWeb, ~100 for FineWeb2). URL filtering (blocklist), language classification (fastText), and quality/repetition filters. FineWeb-Edu uses a classifier trained on Llama-3-70B annotations. Document Fuzzy Deduplication (MinHash) Extensive ablation studies. The final strategy involved deduplicating each Common Crawl snapshot *individually*, as global deduplication was found to discard high-quality data from older snapshots. (penedo2025fineweb2pipelinescale)
Lee et al. (2021) C4, RealNews, LM1B, Wiki-40B-en N/A (focus was on the deduplication technique itself). Substring (Exact) Suffix Array Removed all but one occurrence of any 50-gram (sequence of 50 tokens) that appeared in multiple documents. This method targets boilerplate text. (lee2021deduplicating)
D4 (SemDeDup) Common Crawl (CC-dedup) An initial pass of MinHash-based deduplication was applied before the main technique. Document (Semantic) Semantic Deduplication (Embeddings + K-Means) A multi-stage method that first applies semantic deduplication (SemDeDup) to remove dense clusters of semantically redundant information. (tirumala2023d4)
LSHBloom peS2o (a dataset of 39 million technical documents) N/A (focus was on the deduplication technique itself). Document (Approximate) MinHashLSH with Bloom Filter Index An extension of MinHashLSH that replaces the large, memory-intensive LSH index with a series of lightweight Bloom filters (one for each band). $12\times$ faster, $18\times$ less space usage.